Ato x Cas9FoxP

on Monday, August 31st, 2020 9:16 | by

Knockout of FoxP in Ato.

Print Friendly, PDF & Email

Progress week 26-28

on Monday, July 13th, 2020 1:39 | by

Updates in DTS code

Refreshing dissection skills

Tdc2-Gal4 expression in adult brain
Actin-GFP expression in adult brain

Print Friendly, PDF & Email

Progress for week 25:

on Monday, June 22nd, 2020 1:58 | by

DTS coding
-Added progressbar for data validation
-Updated the progress bar (see figure 1)
-Fixed bug with wrong sample size (see figure 2)
-Fixed bug with unorganized barplots (see figure 2)

Figure 1. Updated progress bar
Figure 2. Correct sample size and restructured barplot


Exp always to the right:
plotOMparams <- plotOMparams[order(plotOMparams$desc),]
plotOMparams$group <- factor(plotOMparams$group, levels=paste(unique(plotOMparams$group)))


Samplesize fix:
samplesizes.annotate(boxes, as.numeric(table(plotOMparams$desc)))

Progressbar:
progress <- c(round(l(100/(length(xml_list)))),round(flycount(100/(totalflies))))

Rescreening:
-Finished rescreening last Thursday. Started to evaluate the new data

Optomotor platform: Ran a few more tests to confirm that the machine was still working, it is. I also adjusted the 0 line so that it is at 0, by readjusting the “zero line” screw. Looks much better now but it is still not perfectly at 0. A difference 0.1 on the computer screen translates to 100 in the evaluation chart.

Optomotor platform:
Ran a few more tests to confirm that the 0 line is always at 0. Readjusted the “zero line” screw. Looks much better now. It is still not perfectly at 0 but a difference of 0.1 in the chart translates to 100 in the evaluation graph.

Print Friendly, PDF & Email

Recent experiments

Print
Print Friendly, PDF & Email

color change

on Monday, May 11th, 2020 11:38 | by

Print Friendly, PDF & Email

FoxP cond

on Monday, April 20th, 2020 2:43 | by

Print Friendly, PDF & Email

New problems with the set-up

on Wednesday, April 15th, 2020 8:46 | by

The measuring device needed repairing, after aperently some glued part got loosloosened and produced no signal anymore. The divice is now producing a signal again, but it is shifted.

The direction of the fly seems to be the main thing. Same fly faceing to the laser (left) and away (right)

Print Friendly, PDF & Email

Updates on optomotor response software

on Friday, March 20th, 2020 3:19 | by

  • Added a timer
  • Previous version generated faulty xml files. This bug has been squashed (see image below)
  • Display window of the software now also prints which direction the striped arena turns
The highlighted red areas are the ones that caused issues. Contingency has also been added to the xml file from the optomotor response (not shown). This did not cause any errors when running the DTS analysis Rscript but I thought it looked nicer having it there.
Print Friendly, PDF & Email

Control cross

on Wednesday, February 19th, 2020 1:20 | by

Control cross of UAS-Cas9;UAS-gFoxP with ELAV-GAL4;UAS-Cas9 is working. Flys are not able to fly and are walking poorly.

Control cross of Hs-Gal4 with UAS-Stinger is still not working.

Print Friendly, PDF & Email

UASCas9;UASgFoxP X ElavG4;TubGal80

Update on coditional FoxP manipulation. Either no manipulation at 18°c or switch to 30°c for adult flys.

Top: switch to 30°c for adult flys.
Bottom: no switch in temperatur
Print Friendly, PDF & Email